The PSL shutter was closed. The beam path blocked two places. High voltage power supplies to IOO and OMC PZT were checked to be off. Oplevs are off.
The south arm green cavity was misaligned yesterday
We would like to keep the vent speed at 1 torr / min. I'm venting with N2 now up to 25 PSI. We have 3 cylinder of instrument grade air inside the lab. Additional supply will arrive later. It can be as late as 1pm